Can you help me convert this to point slope form or standard form? y-3=-2(x=3)
Thanks

Found 2 solutions by stanbon, oberobic:
Answer by stanbon(75887)

convert this to point slope form or standard form?
y-3=-2(x+3)
--
y = -2x-6 + 3
--
y = -2x - 3
---
slope = -2
intercept = -3
===================
Cheers,
Stan H.

Answer by oberobic(2304)

I assume you mean: y-3 = -2(x+3)
.
Slope-Intercept Form: y = mx + b
.
Muliply through to remove the parentheses:
y-3 = -2x -6
.
Add 3 to both sides:
y = -2x -3
.
Standard Form: Ax + By = C
y-3 = -2(x+3)
y-3 = -2x -6
.
Add 2x to both sides:
2x + y -3 = -6
.
Add 3 to both sides:
2x + y = -3